### v1

# Les Trente Glorieuses: 29 years of helioseismic observations with the Luminosity Oscillations Imager

Abstract
Context. The Luminosity Oscillations Imager (LOI) of the Variability of Solar Irradiance and Gravity Oscillations (VIRGO) instrument aboard the SoHO mission has been operating for almost the past 30 years.
Aims. The goal of this article is to report on the observation of p-mode parameters throughout these 30 years that covers two solar cycles using the LOI scientific and guiding pixels as well as the guiding high-voltage housekeeping. The article also provides a definitive description of the data for the final SoHO archive. The LOI time series analyzed here starts on 1 April 1996 0:00 TAI and ends on 31 March 2025 23:59 TAI with a 60-s cadence.
Methods. I use the level-0 data of the LOI for computing level-2 corrected time series for each pixel taking into account engineering calibration, presence of attractors (locked values), the SoHO-Sun distance, the spacecraft roll I use the scientific pixels for deriving Sun-as-a-star signals, individual signals, East-West and North-South difference signals, guiding pixel signals and East-West and North-South high voltage signals to compute 29 one-year power spectra. A 29-year long time series of the Sun-as-a-star signal is also used for pushing the detection of p modes to low frequency. The power spectra are globally fitted (fully over the p-mode envelope) using Maximum Likelihood Estimators using a C++ code with two versions of the code depending on the presence of modes (progFIT and guiFIT). The Fourier spectra for the individual signals are locally fitted (around each p-mode) using Maximum Likelihood Estimators using an IDL code for .
Results. I report on the effect of solar activity upon mode frequencies, linewidths, height and energy rate. I report on the variation as a function of frequency for frequency, coefficient and linewidth changes, as well as the average over the degree and the frequency of these changes. Using the 29-year time series, I report on the frequencies, linewidths and mode height fitted with progFIT. Using the collapsogram technique, I also report on the detection of modes below 1600 Hz, making the lowest frequencies detected with an instrument observing the Sun in intensity. I also report on the detection of p mode in the high voltage and guiding pixel signals with a mode height about 5 to 10 times larger than what is observed in the Sun-as-a-star signal for . The ratios of the observed mode visibilities for the different signals are provided following a calibration of the size of the guiding pixels. While the visibility ratios for the signals excluding the limb are in good agreement with theory, those covering the solar limb are in strong disagreement.
Key Words.:
Sun – solar – p modes1 Introduction
Helioseismology started in the late 1970’s with the first detection of global solar p modes by Claverie et al. (1979), confirmed by Grec et al. (1980), using radial velocity measurements; and by Woodard & Hudson (1983) using irradiance measurements with ACRIM111Active Cavity Radiometer Irradiance Monitor (Willson, 1979) onboard SMM222Solar Maximum Mission. These breakthroughs were followed by the first inversion of the solar structure performed by Gough (1984) and Duvall et al. (1984) using the so-called Duvall’s law (Duvall & Harvey, 1984). Helioseismology really became a useful tool when it demonstrated that the equation of state could be in error, and was indeed in error (Kosovichev et al., 1992).
The diagnostic potential of helioseismology for our physic had already been perceived in the mid 80’s when the space mission SoHO333Solar and Heliospheric Observatory, an ESA/NASA mission (Special volumes of Solar Physics, 1995, 162;1997, 170; et 1997,175) was approved. The need for uninterrupted data can be fulfilled by a space mission, but also by ground-based networks of instrument, networks on which the Sun never sets: GONG444Global Oscillation Network Group (dedicated volume of Science, 1996, 272), IRIS555Investigation on the Rotation and Interior of the Sun (Egamberdiev & Fossat, 1991) and BiSON666Birmingham Solar-Oscillations Network (Elsworth et al., 1991).
The detection of solar oscillations using solar radial velocities is feasible from the ground. It is not so obvious how to perform such measurements using irradiance or radiance fluctuations because of transmission variations through the Earth atmosphere. Following the detection of p modes with the ACRIM instrument, Fröhlich & Van der Raay (1984) used photometers aboard balloons to reduce the impact of the Earth atmosphere on the measurement. A similar photometer (IPHIR777InterPlanetary Helioseismology by IRradiance measurements, (Fröhlich et al., 1988)) was flown aboard the PHOBOS II888Soviet Mars exploration mission, (See Siddiqi, 2016) mission. This instrument returned excellent helioseismic measurements as shown by Toutain & Fröhlich (1992).
In 1987, the SoHO mission was selected by the European Space Agency as part of it Horizon 2 programme. The mission consisted of in-situ, coronal, particle, and helioseismic instruments. Following the at-the-times recent developments in helioseismology, three helioseismic instruments were included for obtaining solar radial-velocity and irradiance measurements using either imagers or looking at the Sun as a star. GOLF999Global Oscillations at Low Frequency (Gabriel et al., 1995) uses a Sodium cell to measure solar radial velocities look at the Sun as a star, while MDI 101010Michelson Doppler Imager (Scherrer et al., 1995) uses a one-million pixel detector for obtaining spatially resolved data. The VIRGO111111Variability of IRradiance and Gravity Oscillations (Fröhlich et al., 1995) instrument measures the irradiance of the Sun as a star, the spectral radiance with photometers, and the spectral radiance spatially resolved over the solar disk. The latter is performed by the Luminosity Oscillations Imager (LOI) whose concept was described by Andersen et al. (1988), and later built as a space instrument (Appourchaux et al., 1997).
The LOI is one of a kind, that is the only instrument that detected in intensity the low degree p modes from the ground (Appourchaux et al., 1995), and that has been measuring the low degree p modes from space since 1996. Its design was specifically aiming at detecting modes with a different sensitivity for degree greater than 1; thereby optimizing the chance of detecting the solar g modes using different signals. Instruments that detects p modes in velocity, such as GOLF and BiSON (Broomhall et al., 2007) or HMI121212Helioseismic and Magnetic Imager (Scherrer et al., 2012) and GONG (Appourchaux, 2024), benefit from a better signal-to-noise ratio allowing them to detect modes below 1 Hz. In addition, instruments such as GONG, MDI and HMI make images of the Sun permitting them to derive by inversion the internal structure from the surface below the convection zone. Nevertheless, it must be emphasized that scientific progress should not always be ranked on the ability of making interesting discoveries, but also on the ability of confirming new discoveries. The failure to confirm g-mode detection can serve as an example (Appourchaux et al., 2000, 2010).
The LOI is not just merely another helioseismic instrument, it can contribute to the understanding of several sources of systematic errors. For instance, it was demonstrated by Toutain et al. (1997) that observations of solar p modes in intensity look indeed different from those performed in velocity. This resulted in the discovery of p-mode profile asymmetries (Toutain et al., 1998). It was also shown by Schou et al. (2002) that rotation inversion are very sensitive to different frequency inferences made with different instrument. Last but not least, the LOI has a different mode sensitivities as it observes in intensity; the mode leakage is therefore significantly different from that of velocity instruments, as are the systematic errors associated with this leakage.
The goals of this article are: 1) to explain how the time series of the LOI pixels are constructed, to be provided to the solar community; 2) to show that an instrument detecting solar p modes in intensity confirms most of the discovery performed in velocity; 3) to report on the limb amplification of the p modes. The LOI time series analyzed here starts on 1 April 1996 0:00 TAI and ends on 31 March 2025 23:59 TAI with a 60-s cadence.
The first section of the article describes the data reduction, and production of level-1 data. The second section explains which times series were computed. The third section explains how the resulting power spectra were fitted for p-mode parameter extraction. The following sections provide results of the variations of the p-mode parameters over two solar cycles, the impact of gaps, the frequency error bars, the mode amplitude and height. The last section gives results from the various guiding signals before concluding.
2 From level 0 to level 1: calibration
The design of the LOI instrument is described in Appourchaux et al. (1997). The data reduction has been heavily modified since then. Especially, the occurrence of locked value on the pixel data, known as attractors, requires a special treatment for identifying them and flagging them. The attractors are values that are locked at some binary values, locking that is most likely occurring in the Voltage-to-Frequency converters used for the Analog to Digital conversion in the VIRGO instrument. The Sunphotometers (SPM) of VIRGO are also affected by this locking. Other corrections related to the loss of the pointing of the High Gain Antenna (HGA) are also required. All these steps are described hereafter.
2.1 Level 0: depacketization
After depacketizing of the transmitted data, the level-0 data are built from the three transmissions of the VIRGO data; this is done at the VIRGO data center (VDC) located at the Instituto de Astrofisica de Canarias, Tenerife, Spain. Preliminary detection of the attractors is done at the VDC but I use my own algorithm described hereafter for flagging them.
2.2 Level 0: attractors detection and flagging
The attractors are best detected in the raw binary data before conversion to engineering values. The detection is performed day by day, pixel by pixel. The procedure is as follows:
First, detect if there are any jumps greater than 0.4% of the median in consecutive data points
Second, if no jumps are detected proceed with the next step, otherwise no attractor detection performed.
Third, I compute the mean and the rms of the data. We arbitrarily limit the detection of the attractors to values within of the mean.
Third, I compute the histogram of the data using a bin size of 1% of the data variation (maximum-mininimum)
Fourth, the histogram is binned with a box car over 30 bins (with edge mirrored)
Fifth, I compute the ratio of the histogram to that of the binned histogram.
Sixth, the attractors are then detected when the ratio is greater than 5.6, where the is the estimated value of the rms ratio for the associated bin value (shot noise). The attractor values are then replaced by a value of -10.
The procedure works in nearly all circumstances. Given the fact, that this procedure is applied daily more than 120, times in total in the course of 29 years, there are still circumstances that would make this procedure fails. For instance, when there is a sunspot appearing on a pixel there is a large variation of several % that would distort the histogram. In order to reduce, the over detection or under detection, we decided to conduct a visual check of the detected attractors, especially when the daily variations are larger than some predetermined value (¿0.3%) typical of a small sunspot crossing a pixel.
Figure 1 gives an example of attractor detection. It shows the first detected attractor. The occurrence of these attractors does not depend upon the pixels but on the digital value itself. Since the instrument has a decreasing throughput of about a factor 16 in 29 years, when the pixels with the smallest flux (Pixels 9 to 12) have an attractor, the other pixels with larger surface will also have the attractor at the same value at a later time. Figure 2 gives an histogram of all of the digital values corresponding to attractors. There is a clear periodicity in the inverse of the binary value shown by Fig 2. The periodicity corresponds to a ratio between the inverse binary value of about a factor 1.0625, which exactly corresponds to a multiplication by 1+1/16. In binary calculation, this multiplication corresponds to adding the original binary value (1) to the binary value shifted down by four bits (1/16). The fact that there is also a correspondence with a factor two which is a binary value shifted down one bit may give clues to the source of the attractors. There are also counters behind the VFC that can count up to 24 bits. It is possible that the locking occurs not in the VFC but more likely in the counters themselves.
The flag of the attractors are kept until the outliers are cleaned (See Section 3.2). At most, there are no more than six attractors at any one time, thereby affecting half of the LOI pixels. The impact of the attractors on the data is studied in Section 3.2.
2.3 Level 1: Engineering data conversion
The level-0 data are then calibrated to engineering units using the calibration performed on the ground before launch. The calibration of the offset for each pixel was refined in flight with the calibration-mode switch off. The data are then corrected for temperature effects and pixel-sensitivity variations.
2.4 Level 1: Sun distance and spacecraft attitude corrections
The 16 pixels are then converted to flux measured at 1 AU using the SoHO-Sun distance derived from the orbital data (See Appourchaux et al., 1997, for the configuration of the pixels on the detector). The correction takes into account the real shape of the pixels, the limb darkening at 500 nm, and the in-flight size of the solar image at 1 AU. The precision of the correction is better than 0.1 % for all of the scientific pixels. In addition a relativistic correction is applied taking into account the spacecraft velocity with respect to the Sun; this is minute. All of these corrections are described in the LOI VDC document available on the VIRGO home page under www.ias.u-psud.fr/virgo/html/software.html. In addition, for taking into account the pixel response that depends on the the spacecraft (S/C) attitude, I used the sensitivities derived by Appourchaux et al. (1997) for correcting for the impact of the varying angle of incidence of the sunlight on the LOI instrument.
3 From level 1 to level 2
3.1 Roll maneuver correction
After May 25, 2003, the HGA failed to move for sending data towards the Earth. It was then decided to leave the HGA in a fixed position and rotate the spacecraft by 180 degrees every 3 months for transmitting to Earth131313See https://soho.nascom.nasa.gov/soc/soho_events/SOHO-Spacecraft-Events.csv. When the S/C is rolled by 180 degrees, the pixel numbering is redefined in the process, i.e. pixels 1, 2, 5, 6, 7, 8 becoming pixels 4, 3, 8, 7, 6, 5 and vice versa; and the guiding pixels 13, 14 becoming 16, 15 and vice versa.
During the roll maneuvers, the rotation is not done exactly around the line of sight of the instrument requiring an adjustment the pointing of the instrument. In the process, this slight depointing produces a slight change in the flux measured by each pixel. The impact of the roll is to slightly change the pixel response. The correction of these changes are done for each pixel, as follows:
First, get date and time of the roll by using the S/C attitude data
Second, the pixel flux is fitted over one day ending 200 minutes before the roll, and starting 200 minutes after the roll.
Third, the flux before and after the roll are then extrapolated for each fit at the time of the roll.
Fourth, the correction is done by normalizing the flux after the roll to that of the flux before the roll.
This normalization makes sure that the time series is homogenous since the start of the mission.
3.2 Outliers cleaning
The various interruptions due either to the extended loss of contact (the so-called SoHO vacations, mid 1998), the roll maneuvers, the presence of attractors or other unforeseen events produce large jumps in the pixel data that need to be corrected for. We use an automatic outlier detection based upon the Peirce criterion (Peirce, 1852; Gould, 1855). The correction for each pixel is done as follows:
First, detect outliers with the Peirce criterion over a duration of 10 days. The detection and correction are done automatically. The detected outliers are filled by interpolation.
Second, detect outliers with the Peirce criterion over a duration of 360 days. The detection can be done either automatically (default) or interactively (choice of the user). If the correction is done interactively, the user chooses two locations over which to perform the interpolation. The latter is used when very large jumps cannot be corrected automatically. This step is done recursively until no correction is necessary.
Third, detect outliers with the Peirce criterion over a duration of 40 days with the same procedure as for the previous step
During this process, the attractors are detected as outliers and corrected by interpolation, thereby filling the gaps created by these attractors. Figure 3 shows the impact of the attractors on the duty cycle of the data. At the end of the mission, the data have attractors about 40% of the time. The increase of the number of attractors is mainly due to the fact that the photon flux on the pixels decreased by factor 16 over the 29 years, hence the change in digital count is smaller at the end of the 29 years compared to the beginning, making the attractors lasting longer at the end than at the beginning. But since the attractors are not always continuous, the interpolation can easily fill them in when they are shorter than five minutes. The number of pixels having an attractor lasting more than five minutes is at most 5%, while it is less than 0.5% for two attractors.
3.3 Detrending and filtering of pixel data
For each pixel, the data are detrended using a triangular two-day smoothing window. The relative values are then computed and all spikes greater than 30 are removed. Since the 12 pixels may have some missing value (set to zero), a further filtering is done by setting all pixels to zero if at least one pixel is missing in all the three quadruplets: Pixel 1 to 4, Pixel 4 to 8, and Pixel 9 to 12. The 12 pixels are then not set to zero if for any datum at most two pixel quadruplets have missing data. The time series analyzed of all 16 pixels are available from soho.nascom.nasa.gov/data/archive.html.
4 Generation of time series
4.1 Sun-as-a-star time series
The data set for the Sun-as-a-star time series or Full-disk (FD) time series is derived from the average of the 12 level-2 scientific pixels. The time series are made of two sets: one time series of 29 years duration, 29 time series of one-year duration. Examples for the power spectra for the first year and the last year are shown on Figure 4. The power spectra are corrected by dividing by the duty cycle (See Figure 3).
4.2 Spherical harmonics extraction
For combining the pixel data, I used the spherical harmonic decomposition computed as in Appourchaux & Andersen (1990), using the limb darkening from Allen (1975). The filters are normalized according to Eq. (17) of Appourchaux et al. (1998a). This normalization allows a symmetrical leakage matrix (Appourchaux et al., 1998a). Example for such leakage matrix can be found in Appourchaux et al. (1998c). Since the apparent solar diameter changes along the orbit, the spherical harmonic filters vary slowly with time. To alleviate this problem, time-independent filters are applied to the pixels for extracting the modes. The filters are computed as an average over one year of filters computed weekly. Over this period the angle has a mean close to zero. Gizon et al. (1998) showed that it was possible to detect the influence of a non-zero angle on the p-mode data. Therefore, this average minimizes the effect of the angle on the data. After combining the pixels with the filters, the time series are Fourier transformed; the positive frequencies provide the signal for , and the negative frequencies provide the conjugate of the signal for Appourchaux et al. (1998a).
For each harmonic, I extracted two data sets: one time series of 29 years duration and 29 time series of one-year duration.
4.3 East-West and North-South time series
I also extracted the difference between the North-South (NS) pixels and the East-West (EW) pixels. For the NS pixels, this is simply the same data set as for . For the EW pixels, this is the difference between the complex time series and its conjugate. For each, I extracted 29 time series of one year duration. These data sets were used to compare the result obtained with the high voltage (HV) signals, which are sensitive to the difference between the North and South guiding pixels and between the East and West guiding pixels.
4.4 EW / NS HV and guiding pixel time series
I also extracted from the engineering data, the high voltage time series for the Y direction (North-South) and the Z direction (East-West). The voltages were converted to arcsec then to ppm. The value of 8.8 ppm/marcsec is derived from the outer radius of the scientific pixels given above and from the calibration performed in April 1996 (Appourchaux et al., 1997). Figure 5 shows two power spectra at the beginning of the mission and near the end of the mission. The improvement in signal-to-noise ratio is striking but is not understood. On the other hand, the guiding pixels also show the p modes but only for the first year (1996) (Figure 6) since the noise increases by such a factor that no modes can be detected after 1997, which is not understood either. Similar power spectra were derived with six months of data by Appourchaux & Toutain (1998).
5 p-mode parameter extraction
They are 30 Sun-as-a-star power spectra, 29 North-South difference power spectra, 29 East-West difference power spectra, two guiding pixel power spectra, 29 North-South high voltage power spectra, 29 East-West high voltage power spectra, and 29 15 power spectra.
5.1 Power spectra fitting
The power spectra are fitted using Maximum Likelihood Estimation (MLE) (Schou, 1992; Appourchaux et al., 1998a). The power spectra are fitted globally (fully over the p-mode envelope), an approach first used for fitting stellar power spectra as in Appourchaux et al. (2008). In that latter article, the MLE for HD49933 was made difficult due to the large linewidth of the modes making the identification of the pair with respect to the difficult. It was already suggested by Appourchaux et al. (2008) that a Bayesian approach would be better providing a more conservative identification of the modes; such an approach was implemented by Benomar et al. (2009). This Bayesian approach that I suggested is now widely used in asteroseismology (Handberg & Campante, 2011; Corsaro & De Ridder, 2014; Lund et al., 2017). The Bayesian fit of the power spectra is remarkably efficient when the S/N ratio in the power spectrum is low, less than about 3, but when the signal ratio is higher the MLE fit and the Bayesian fit provide to the same estimators with very similar error bars. In addition, the Bayesian fit requires the creation of Monte Carlo Markov Chains which can be very slow in any interpreted language (Python, IDL).
The code used by Appourchaux et al. (2008) for the CoRoT star HD49933 was indeed an IDL code. There were already attempts to write a C++ code (progFIT) based on this IDL code (Neiner & Appourchaux, 2004) which was not yet available in 2008. The development of that code continued until 2010, but the code was not still fully functioning (Neiner, private communication). The code was then picked up from its ashes in 2021 when its use for the Planetary transits and oscillations of stars (PLATO) mission was considered an asset given the large sample of solar-like stars (About more than 20, stars for the core P1 sample, Rauer et al., 2025). The code was then made fully functional and was tested for the PLATO mission on hundreds of solar-like stars. The code can be found on Gitlab141414git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
/msap3_04/progfit. The specificity of progFIT is that it uses the Minuit2 library for finding the highest likelihood, but it not only finds the highest likelihood (or (likelihood) minimum) it also check that the distance between the minimum found and the extrapolated minimum is small compared to a given criteria, i.e. providing the estimated distance to minimum (EDM). The Hessian (used for getting the parameter errors) is computed on the fly as well as the EDM. The EDM is a representation that in the found minimum is within a typical distance in the hypersphere of the parameters modeling the power spectrum. The EDM is a dimensionless quantity that can be written as
where is the relative distance to minimum in units of , and is the mean relative distance over the parameters. For 200 parameters and a mean of 3%, the resulting EDM is 0.1. I used this latter value for the convergence of all fits. For an EDM of 0.1, when the number of parameters is decreased to 100, the mean distance to the minimum increases by a factor , still sufficient for most applications.
For progFIT, the typical speed is 22 minutes on an Apple M3 Max, for a one-year duration time series, with 19 quadruplets =0, 1, 2 and 3 for a total of 166 parameters: 10 parameters for the background (three Harvey-like profiles with exponent of 4 and one white noise), one common splitting for all degrees, one common inclination angle for all degrees, one common ratio between the heights of the =2 modes to =0 modes, one common ratio between the heights of the =3 modes to =1 modes and 8 parameters per quadruplet (four frequencies, two linewidths (one for the pair and one for the pair), two mode heights) assuming a symmetrical Lorentzian profile; the ratio within a multiplet follows Toutain & Gouttebroze (1993). For the 29-year duration time series, it takes 8 hours with 23 quadruplets.
Following the successful development of progFIT, another code was developed for fitting the NS and EW difference signals, and HV guiding signals. Since these signals are by construction not sensitive to =0 modes, the new code (guiFIT) took into account this lack of visibility, but also the different visibilities in a multiplet. The code can also be found on Gitlab151515git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
/msap3_04/guifit.
The progFIT code was used 62 times on the FD and guiding pixel power spectra, while guiFIT was used 116 times. In total, the two codes were used less than about 180 times resulting in a bit less of 3 days of computing on a single processor, in comparison the former IDL code would have taken about 150 days of computing. With several cores such as the M3 Max, the duration reduces even further to about five hours.
5.2 Fourier spectrum fitting
This technique, known as phase fit, was introduced by Schou (1992) and later refined by Appourchaux et al. (1998a). The method relies on the knowledge of the leakage matrix for inferring the p-mode parameters. This method can only be applied when the leakage matrix for a given degree can be inverted (Appourchaux et al., 1998a).
The p-mode parameters are usually locally extracted (around each p-mode) from the Fourier spectra using MLE (Schou, 1992; Appourchaux et al., 1998a) with an IDL code. The splittings were derived from Clebsh-Gordan coefficients as in Ritzwoller & Lavely (1991), the decomposition is given in Appendix A. Three pixel noises were used for modeling the solar noise, while the leakage and noise covariance matrix were derived according to Appourchaux et al. (1998a). Examples of Fourier spectra for various and can be found in Appourchaux et al. (1998c).
Here I did not study the odd -coefficients that are related to the solar differential rotation. Although results on these coefficients are available with the LOI (Appourchaux et al., 1998b) their use has to be combined with higher degree modes for getting a proper rotation inversion (See Appourchaux et al., 2002). Since the LOI does not provide high degree modes, work using the GONG and HMI / MDI data provides odd -coefficients covering all degrees from 0 to 200 with a single instrument (Korzennik, 2023) allowing homogenous rotation inversion.
6 p-mode parameters as a function of solar activity
The dependence of the p-mode parameters upon solar activity was studied against the flux at 10.7 cm. This flux is the manifestation of the existence of magnetic fields on the surface of the Sun.
The 10.7-cm flux could be downloaded at the SIDC (Solar Influence Data Analysis Center) in 2023 with a time series running from April 1st 1996 to March 31st, 2021; the data are at the time of writing no longer available. The flux can now be found daily thanks to Space Weather Canada161616spaceweather.gc.ca/forecast-prevision/solar-solaire/solarflux/sx-5-en.php, and the two time series were merged into a single time series from April 1st 1996 to March 31st, 2025, for producing a yearly average. Since I used the 29 yearly time series of the different signals for that study, I also averaged the flux over the same period of one year. The 29 power spectra were fitted using a variable and a fixed noise background. The two fits were used to reduce the bias on linewidths and mode heights that occurs when one uses a different background (See for instance Appourchaux et al., 2014).
6.1 Frequencies
The impact of solar activity on the frequency and the various even -coefficients of the splitting was first studied by Libbrecht & Woodard (1990). The even coefficients are directly related to the decomposition of the magnetic field perturbation on even-order Legendre polynomials (Antia et al., 2001).
For studying the frequency variation over the solar cycle, I used Eq. (5) of Chaplin et al. (2004) that includes a correction for the mode inertia and a function that measures the variation of the frequency shift over frequency, normalized to 1 at 3 Hz. This latter function was fitted from Figure 1 of Chaplin et al. (2004) as:
where is the frequency normalized to 3 Hz. The weighting as provided by their Eq. (5) is a way to relate all frequency variations to a single frequency of 3 Hz, then can provide the frequency shift at any given frequency if needed. The weighting scheme is used over the frequency range from 2400 Hz to 3500 Hz for obtaining the variations for each degree, and over all degrees up to .
6.1.1 Dependence on the degree and frequency
Figure 7 shows the frequency shift for =0, 1, 2 and 3, either fitted globally for the Sun-as-a-star signal and the difference signals of the resolved image (EW and NS) or locally with the resolved image. The maximum frequency variation is typically 0.3 to 0.4 Hz, depending on the degree. For the mode, there is no difference between the local and the global fit. For the other degrees, there are differences due to the impact of the - and -coefficients upon the frequency shift. For different from 0, each signal has a different dependence upon the - and -coefficients. The frequency shift for each signal depends upon the visibility of each component. For the NS differences and the HV Z, the signals are primarily sensitive to the modes for . For the EW differences and the HV Y, the signals are primarily sensitive to the modes for . For the FD signals, the sensitivity varies with the degree for which the dependence of the measured frequency was given by Appourchaux & Chaplin (2007). Table 1 gives the frequency dependence for all these signals and degrees. Figure 7 clearly shows the result of Table 1.
Using the resolved data, I derived the dependence of the -coefficient upon the 10.7-cm flux for the three different degrees. Figure 8 shows the slope of the linear fit of to the 10.7-cm flux. Over the 29 years, the maximum variation of the -coefficient is about 0.1 Hz (with the impact of the being negligible) resulting in a difference between the EW- and NS-signal frequency shift of about 0.3 Hz at the maximum of activity (See Fig. 7 and Table 1).
| Signal | |||
| EW / HV Y | - | ||
| Full disk | |||
| NS / HV Z | - |
6.1.2 Degree and frequency averaging
The averaging done over the frequency range can also be combined for averaging over the degrees. Figure 9 shows the averaged frequency shift over various degrees for the FD signals, the NS signals (difference and HV Z) and the EW signals (difference and HVY). It is clear that the difference in frequency shift is due to the different mode sensitivity whose variation with results in a different sensitivity to the solar activity belt. In addition, the variations are the same for the NS / EW differences and the HV Z / HV Y signals (See Table 1).
The averaging over frequency and degree can also be done using a time-travel technique (Vasilyev & Gizon, 2024). Although their claims of being a simpler method than mode frequency fitting is matter of debate, they still find similar variations with solar activity for the mean mode time travel.
6.1.3 Frequency error bars
I also compared the formal error bars obtained with the global MLE estimation with that of the rms deviation of the individual frequencies. Since the frequencies vary with the solar activity, I compensated the variations for each degree and the solar-activity level as given by the 10.7-cm flux. After subtraction of the frequency shift which increases with frequency until 4 Hz (See Figure 10), I was able to infer the rms variations for each degree and frequency. Figure 11 gives the rms and formal errors for each degree for the FD signal, while Figure 12 gives the ratio of the rms errors to the formal errors for =0 to 3. The formal errors are typically smaller by 20% to 30% smaller than the rms errors. Such underestimation was also measured for the 45-year of BiSON splittings (Howe et al., 2023). This underestimation does not contradict the fact that the MLE errors are bounded on the low side by the Cramer-Rao theorem (Appourchaux et al., 1998a). The same bounding occurs with a Bayesian fit since in that case one must use the Van Trees (Van Trees, 1968) inequality taking also into account the information prior. The two inequalities are such that the Bayesian bound is lower than the classical MLE bound (See also Echeverria et al., 2016, for a hands-on application). When there is no information, the two bounds are identical.
6.2 Linewidth and mode heights
The variation of low-degree-mode linewidths and mode heights were reported by Chaplin et al. (2000) using BiSON data. Similar variations were found by Jiménez-Reyes et al. (2003) for GOLF and VIRGO data. Kiefer et al. (2018) reported on the variation of these mode parameters for the GONG instruments over two solar cycles. The increase in mode linewidth is related to the decreasing size of granulation that drives the excitation mechanism of the modes (Kiefer et al., 2018). Such variations in granular size with solar activity were indeed measured by Ballot et al. (2021) but with a one-year delay. Changes in linewidths using VIRGO/SPM data were also reported by Toutain & Kosovichev (2005) during half a solar cycle. Hereafter, I report on variations over two solar cycles taking into account the varying duty cycle in the data.
6.2.1 Impact of gaps
When measuring linewidths and mode heights, one has to take care about how the power is redistributed in the power spectrum. When there are gaps in the data, the frequency bins become correlated (Gabriel, 1994). The influence of such gaps in the power spectrum can be corrected by properly taking account the statistics as shown by Stahn & Gizon (2008). In practice, the power spectrum fitting is done assuming that the frequency bins are not correlated resulting in biasing the mode linewidth as shown by Komm et al. (2000a) for the GONG data, or by Chaplin et al. (2000) for the BiSON data. For studying the impact of gaps on the mode linewidth and mode height, I created 29 times series using the first year as a reference, then applied the data gap structure of each time series to the reference time series, thereby creating 28 new time series. I averaged the mode linewidth and mode height over the frequency range from 2400 Hz to 3500 H (as for frequency), and over the degrees. I did a straight average between over the frequency range, not using the weighting scheme of Chaplin et al. (2004). Figure 13 provides the variation of the mean mode linewidth and mean mode height as a function of the gap fraction, together with a linear fit of the dependence. On average for any 1% additional gap, the mode linewidth is increased by 0.85 %, while the mode height is reduced by 2.1 %. Values of the same order of magnitude were found by Keith-Hardy et al. (2019) for the GONG data. The impact of gap on linewidth and height is the direct result of the dilution of power due the widening of the window function (Gabriel, 1994). The dependence is used for compensating the effect of gaps on the mean mode linewidth and height (See Section 6.2.3).
6.2.2 Linewidth dependence on the degree and frequency
Using the resolved data, I inferred the dependence of the mode linewidth on the 10.7-cm flux as a function of frequency. Then I computed a linear fit of the mode linewidth variation to the 10.7-cm flux. Figure 14 shows the slope as a function of frequency for =1 to 3. It confirms that the mode linewidth dependence to activity is different from that of the frequency dependence (See Fig. 10), confirming what Kiefer et al. (2018) found with the GONG data.
6.2.3 Degree and frequency averaging
I averaged the mode linewidth and mode height over the frequency range [2400, 3500] Hz, and over the degrees. I did a straight average between over the frequency range, not using the frequency weighting scheme of Chaplin et al. (2004)the var. Figure 15 shows the variation obtained over the solar cycle for mode linewidth, height, power and energy rate. In Figure 15, I also show the variations of the linewidth measured with a fixed background but corrected for the effect of biases due to the different background and splitting; the procedure is the same as followed by Appourchaux et al. (2014). One can notice that the biases are indeed corrected by the procedure and that the variations follow the same trend with solar activity. The linewidth variations are correlated with the activity while the mode height and power are anti-correlated. The energy rate is independent of solar activity. The same results were observed by Broomhall et al. (2015) and Kiefer et al. (2018) with the BiSON and GONG data, respectively.
7 Results for the 29-year time series
I used the single time series of 29 years of the FD data for getting robust estimate of mode frequencies, linewidths and heights. Figure 16 shows the power spectrum obtained with the 29 years of FD data together with the ratio of the power spectrum to the fitted model. The regular repetition of the modes can be used to derive an échelle diagram (Grec, 1981). Figure 17 shows the échelle diagram of power spectrum obtained with the 29 years of FD data together with the échelle diagram of the ratio of the power spectrum to the fitted model. The fitted model was obtained with the (progFIT software. Although that software is efficient at recovering nearly all the modes, modes with low signal-to-noise ratio and narrow linewidths are not fully recovered. A Bayesian approach with the known location of these low frequency modes (derived from velocity instrument such as GOLF) is more efficient. In addition, it has been demonstrated that there is an optimal way to smooth the power spectrum for increasing the detectability of low frequency modes (Appourchaux, 2004).
Combining the collapsogram technique (Appourchaux, 2024; Salabert et al., 2009) and the smoothing of the power spectrum, I could detect modes below 1800 Hz (See Figure 18). I used the collapsogram technique together with an optimization of the -, - and -splitting coefficients. The optimization is performed with a genetic algorithm already used by Appourchaux (2020) taking into account the statistics of the maximum (See Appourchaux, 2024). After the optimization, the resulting collapsed power spectrum is fitted using MLE. As shown by Appourchaux (2003), the fitting of averaged power spectrum can be performed using the MLE technique for two degrees of freedom (d.o.f), but scaling of the error bars is required to take into account the number of d.o.f. Table 16 gives the frequencies obtained by globally fitting the power spectrum of the 29-year time series, while Table 17 provides the frequencies obtained at low frequency using the collapsogram.
Figure 19 gives the mode linewidth and mode height obtained using the global fit and the collapsogram (See Tables 18 and 19). Since the averaging is performed over nearly 2 solar cycles with varying linewidths and frequency shifting, the mode linewidths are slightly higher than at low activity. The impact of the frequency shift upon the linewidth was studied by Chaplin et al. (2008) but not taking into account the variation of the linewidth; the typical overestimation they obtained over a solar cycle for the linewidth were less than 3%. For taking into account the change of all parameters over the solar cycle, I simulated the impact of the variations using the measured mode frequency, linewidth and height variations. I created 29 yearly time series for a single p mode to obtain a time series spanning the two solar cycles. Each time series has an injected frequency shift according to Fig 10 and a linewidth and height changes according to Fig 14, all scaled to the 10.7-cm flux. I fitted the resulting power spectrum for four typical frequencies: 2500 Hz, 3000 Hz, 3500 Hz, 4000 Hz. While the resulting fitted frequency shift is the same as the mean injected frequency shifts (averaged over the 29 time series), the mode linewidth variations are about 30% larger, i.e. the linewidth are 5% and 10% larger, at 2500 / 3500 Hz and 3000 Hz; there is no impact at 4000 Hz as there is no resulting mode linewidth variation (See Figure 14). Therefore, the resulting simulated bias on linewidth is only related to the linewidth variation with a typical frequency dependence similar to that of Figure 14. Despite this bias, the mode linewidths are in good agreement with measurements obtained by BiSON (Chaplin et al., 1997; Davies et al., 2014), GOLF (Gelly et al., 2002), GONG (Komm et al., 2000b; Korzennik, 2023), HMI (Korzennik, 2023) and VIRGO/SPM (García et al., 2011). Note that the linewidths measured by Komm et al. (2000b) with GONG are smaller than the other measurements above 3700 Hz.
| Source | radius (NS) | radius (EW) | radius (guiding) | ||
| Allen (1975) | 0.9700 | -0.2200 | 0.941 0.014 | 0.945 0.014 | 0.937 0.006 |
| Neckel & Labs (1994) | 0.9494 | -0.2080 | 0.943 0.014 | 0.947 0.014 | 0.938 0.006 |
8 Results for the guiding signals
Figure 20 shows the mode heights for and for the various guiding signals. It is obvious that the mode heights for the HV signals are about 5 to 10 times larger than those of the FD and resolved signals. On the other hand, the associated linewidths as shown in Figure 21 are very similar, showing that the larger mode heights measured are not a result of a potential bias in the measured linewidths. It was already obvious when comparing Figure 6 with respect to Figure 4 that the mode height in the guiding signals was much than in the FD signals. The higher mode heights shown on Figure 20 were also observed in the MDI instrument by Toner et al. (1999), the HMI instrument and the PICARD mission by Corbard et al. (2013). They all showed that the p-mode power increases towards the limb having power larger by a factor five, for integration over an annulus with a width of 0.5% R⊙. Their results agree with the ratio of mode heights reported.
9 Visibilities and visibility ratios
9.1 Limb darkening and solar radius calibration
The computation of the mode visibilities requires calibrating the limb darkening and the effective size of the solar radius on the LOI detector (See Appendix B). The calibration was done by computing the ratio of the mean of the NS pixels (1, 2. 3, 4) to the Central pixels (9, 10, 11, 12), the ratio of the mean of the EW pixels (5, 6, 7, 8) to the Central pixels, and the ratio of the mean of the guiding pixels (12, 13, 14, 15) as a function of the SoHO-Sun distance in AU. The computation of the relative fluxes was performed using the shape of the pixels and the presence of the tracks as provided by the manufacturer171717Advanced Micro Electronics, Horten, Norway. The tracks are non sensitive part of the detector required for getting the current resulting from the photoelectric effect. The limb darkening is assumed to be expressed as:
where , and is the distance from the Sun disk center expressed in solar radius.
Figure 22 shows the ratios measured during the 29 years by plotting one point per month as a function of the Sun-SoHO distance, together with the fit of these ratios. Unfortunately, this process does not allow to calibrate the limb darkening since there is a 100% correlation between the apparent solar radius and the linear () and quadratic () terms of the limb darkening. Instead, I used two sources for the limb darkening: Allen (1975) and Neckel & Labs (1994). For simplification, the limb darkening of Neckel & Labs (1994) was reduced to that of the two terms above and interpolated at 500 nm. Table 2 gives the radius of the outer dimension of the scientific pixels with respectT to that of the Sun observed at 1 AU derived from the fit shown on Figure 22. This calibration is used for deriving the visibilities for the solar p modes observed either with the scientific pixels or with the guiding pixels. I used the guiding pixel value of 0.937 0.006 (Allen limb darkening).
9.2 Visibilities for the different signals
The theoretical visibilities of the p modes are directly related to the intensity perturbations which are related to the spherical harmonics and the limb darkening (See Appourchaux et al., 1998a, and Appendix B). This is correct to a first approximation but Toutain et al. (1999) showed that this approximation fails at the limb of the Sun because there the photons come from different heights in the solar atmosphere, thereby explaining the enhanced mode heights in the guiding pixels. Further developments were carried by Kostogryz et al. (2021) that included other opacity contributions in addition to that of H-. Developments that are used for understanding the difference between intensity and velocity measurements in HMI (Fournier et al., 2025).
Using the calibrated size of the guiding pixels with respect to the Sun radius at 1 AU, I derived the theoretical visibilities for the FD, the resolved, the EW and NS difference signals, the EW and NS HV signals and the EW and NS guiding pixels. Then the theoretical ratios of visibilities are computed.
From the observations, I compute the average mode height for a given frequency over the 29 time series for all signals, as shown in Figure 20. The ratio is then computed only for the modes spanning the frequency range [2500 Hz, 3500 Hz]. The ratios are then averaged over the modes for getting the mean ratio; the error bars were deduced from the rms value divided by the square root of the number of frequencies. Tables 11, 12, 13, 14 and 15 give the comparison between the ratio averaged over the 29 time series and the theoretical ratio at the mean SoHO-Sun distance of 0.9906 AU.
The radius of the outer boundary of the scientific pixels was also fitted in order to minimize the difference between the theoretical and observed FD ratios. The adjusted value of 0.941 0.08 is very close to the value given in the previous section. This coherence gives confidence in the value of the size and shape of the pixels, and in the understanding of the impact of the SoHO-Sun distance in the mode visibilities.
The FD visibility ratios shown in Table 11 are in good agreement for all degrees but . The observed visibility ratio for is 3.7% of the compared to a theoretical value of 2%. This is a large discrepancy that I did not study in detail but could be related to non-linearities due to the varying p angle (the inclination of the solar axis towards the observer). The visibility ratios of Table 11 are different from those of Ballot et al. (2011) mainly because of the presence of tracks in the FD signal of the LOI, not present in true Sun-as-a-star signals (the track size is about 1.6% of the radius of the solar image).
The resolved visibility ratios shown in Table 12 are also in good agreement within a multiplet, while there is a systematic bias of about 19% with respect to the FD visibilities. The EW and NS difference signal visibility ratios shown in Table 13 are in very good agreement showing no discrepancy.
Finally, the visibility ratios for the EW and NS signals (guiding pixels and HV) are given in Table 14 showing very large discrepancies that are also clear in Figure 20. The same discrepancy appears for the visibility ratios between the HV signals and the guiding pixels in Table 15. One can notice that the theoretical visibilities for the EW and NS signals (guiding pixels and HV) for a given degree are identical, as demonstrated in Appendix C, then so are the ratios of visibilities. I discuss the implication of these observed visibilities in the last section.
10 Discussion and conclusion
In this article, I studied 29 years of observation of the LOI instrument aboard SoHO. I corrected the impact of the so-called attractors that would lock the measured intensity flux to several digital values. The flux intensities were also corrected to take into account the roll of SoHO due to the failed mechanism of the HGA. Although the fluxes dropped by a factor 16, the power spectrum of the Sun-as-a-star time series did not drastically change between 1996 and 2004. On the other hand, the power spectrum of the guiding pixels shows the solar p modes only in the 1996 time series. After 1997, the noise is so high that the modes cannot be detected in the guiding pixels anymore. Surprisingly, the HV signals for the EW and NS directions are much less noisy between 1996 and 2024, typically the noise signals became about four times lower making the detection of p modes much easier. Using C++ codes developed for the PLATO mission, I globally fitted the power spectra of the various signals.
I studied the effect of solar activity upon mode frequency, mode linewidths and mode heights for various signals. I confirm previous finding on the dependence of frequency with solar activity which is similar to the surface frequency effect. These results are new, this is the first time that they are published for intensity measurements. The impact of the -coefficients upon the frequency shift of the different signals have an implication upon the measurement of the frequency shift in other stars (Also studied by Benomar et al., 2023). The NS difference or NS HV mimic the observations of star having a rotation axis inclined pole on (inclination angle of 0), while the EW difference or EW HV mimic that of a star observed with the rotation axis in the plane of the sky. The difference between the frequency shift is larger for the EW signals compared to the NS signals. For any other inclination angle the frequency shift will vary between these two extremes. In other words, the level of activity as measured with the frequency shift depends upon the inclination angle. Extreme care must be taken when drawing conclusions on stellar activity with stars of similar age, radius and mass. This may affect stellar activity studies performed with PLATO (Rauer et al., 2025).
I also studied the heights of the modes detected in the guiding pixels and HV signals. The mode heights in the HV guiding signals are about 5-10 times larger than the modes observed in FD for , while it is about 10-20s times larger for (after correction of the FD visibility). Such higher mode heights are in agreement with other observations performed by Toner et al. (1999) and Corbard et al. (2013) with MDI, HMI and the PICARD mission. On the other hand, the visibility ratios for the EW and NS signals (guiding pixels and HV) reported in this article are clearly not in agreement with the expectation of the simpleI theory provided in Appendix C. There are several discrepancies with theory to be noted (See Table D.4 and D.5): 1) the mode visibilities of the EW signals are systematically higher than those of the NS signals; 2) the mode visibilities of the modes are significantly higher than those of the modes for the guiding pixels and the HV signals. If one compares the results of the HV signals with the EW and NS difference (See Table D;3), it is clear that for these latter signals 1) the mode visibilities in EW and NS are identical for and ; and 2) the ratio of mode visibilities of the to the modes are in agreement with theory. It means that the intensity perturbation at the center of the disk () down to follows a product of the limb darkening with a spherical harmonic (Eq. C.1).
When this is not the case, one can use the formalism given by Toutain et al. (1999) for deriving the visibility ratio of with respect to . I show in Appendix C that the ratio of these visibilities for either the EW or NS signals is . For spherical harmonics, this ratio is about 0.2 (See Table D.4). Using the relative intensity perturbation provided by Toutain et al. (1999) in their Figures 5 and 6 for (used by ) and for (used by ), one can deduce that the ratio should be much less than 1 as well. The intensity perturbations provided by Toutain et al. (1999) do not contradict the simple theory provided by Table D.4. The theory of Toutain et al. (1999) was revised by Prokhorov et al. (2014) and Kostogryz et al. (2021). The latter used additional opacity in addition to the sole contribution of bound-free transitions of H- included in the model of Toutain et al. (1999). Unfortunately, Kostogryz et al. (2021) did not provide the intensity perturbations for and . The computation of such perturbation using the methodology of Kostogryz et al. (2021) is beyond the scope of the current article.
This work may also serve future purposes for showing what can be achieved with a low resolution instrument detecting p modes in other stars. Up to now, the observations of stellar p modes by CoROT (Baglin, 2006), Kepler (Gilliland et al., 2010), TESS (Ricker et al., 2014) and the coming PLATO (Rauer et al., 2025) missions only allow the detection of low degree modes below . In the future, we hope to have large interferometers in space such as the concept of the Stellar Imager proposed by Carpenter et al. (2010), which was aiming at providing about 30 resolution elements across a star (). It would require the formation flying of 30 2-m diameter telescopes on a km baseline. Such challenging developments are not impossible as a realistic road map for formation flying for space-based interferometry has been devised (Monnier & al., 2019).
Finally, this work will be used as the reference for the data reduction of the LOI instrument for which data will be archived at the European Space Astronomy Centre (ESAC) in Villafranca, Spain in the SOHO data archive. At the time of writing, the end of the SoHO mission is planned for September 2026 (D. Müller, private communication).
Acknowledgements.
This article benefited from fruitful discussions I had with Bo Andersen, Antonio Jiménez and Torben Leifsen held during a memorable meeting in Svalbard. This article represents the work performed during 3/4 of my career during which I owe a lot to several of my forefathers: Jacques-Emile Blamont who introduced to the world of helioseismology and political maneuvering; Pierre Connes who introduced to the world of well-thought instrumentation, of spectroscopy and exoplanet search; David M. Rust who had more faith and confidence in me that I will ever have which enabled me to be more aware of what I could do; Claus Fröhlich who permitted me to be in charge and to develop the LOI instrument. I owe a lot also to my career-long colleagues: Bo N. Andersen who was at the start of the LOI concept, long standing friend whatever the circumstances; Antonio Jiménez still kicking the VIRGO Data Center and provider of the Level 0 data. I owe also to the many ESTEC engineers that allowed me to design, to develop, to test the LOI especially Udo Telljohann. Last but not least, I could not have done this career without the long time support of my wife, Maryse, there since the university years; my two sons, Kévin and Thibault, and our many gone or still-living cats. A special mention to Roger-Maurice Bonnet, Alan Gabriel and Jean-Claude Vial for permitting me to go back to instrumentation development in France. SoHO is a mission of international collaboration between ESA and NASA. The development of the (progFIT code was initiated by Bernard Leroy and Coralie Neiner at my prodding, code which was transformed into a regularly usable code with the major contribution of Claude Mercier, and the timely contribution of Vincent Buttice.References
- Allen (1975) Allen, C. W. 1975, Astrophysical Quantities (The Athlone Press)
- Andersen et al. (1988) Andersen, B. N., Domingo, V., Jones, A., et al. 1988, in Seismology of the Sun and Sun-like Stars, ed. V.Domingo & E.Rolfe (ESA SP-286, ESA Publications Division, Noordwijk, The Netherlands), 385
- Antia et al. (2001) Antia, H. M., Basu, S., Hill, F., et al. 2001, MNRAS, 327, 1029
- Appourchaux (2003) Appourchaux, T. 2003, A&A, 412, 903
- Appourchaux (2004) Appourchaux, T. 2004, A&A, 428, 1039
- Appourchaux (2020) Appourchaux, T. 2020, A&A, 642, A226
- Appourchaux (2024) Appourchaux, T. 2024, Sol. Phys., 299, 155
- Appourchaux et al. (2002) Appourchaux, T., Andersen, B., & Sekii, T. 2002, in ESA Special Publication, Vol. 508, From Solar Min to Max: Half a Solar Cycle with SOHO, ed. C. Fröhlich and A. Wilson, 47–50
- Appourchaux & Andersen (1990) Appourchaux, T. & Andersen, B. N. 1990, Sol. Phys., 128, 91
- Appourchaux et al. (1997) Appourchaux, T., Andersen, B. N., Fröhlich, C., et al. 1997, Sol. Phys., 170, 27
- Appourchaux et al. (2014) Appourchaux, T., Antia, H. M., Benomar, O., et al. 2014, A&A, 566, A20
- Appourchaux et al. (2010) Appourchaux, T., Belkacem, K., Broomhall, A.-M., et al. 2010, A&A Rev., 18, 197
- Appourchaux & Chaplin (2007) Appourchaux, T. & Chaplin, W. J. 2007, A&A, 469, 1151
- Appourchaux et al. (2000) Appourchaux, T., Fröhlich, C., Andersen, B. N., et al. 2000, ApJ, 538, 401
- Appourchaux et al. (1998a) Appourchaux, T., Gizon, L., & Rabello-Soares, M. C. 1998a, A&AS, 132, 107
- Appourchaux et al. (2007) Appourchaux, T., Leibacher, J., & Boumier, P. 2007, A&A, 463, 1211
- Appourchaux et al. (2008) Appourchaux, T., Michel, E., Auvergne, M., et al. 2008, A&A, 488, 705
- Appourchaux et al. (1998b) Appourchaux, T., Rabello Soares, M. C., & Gizon, L. 1998b, in IAU Symposium, Vol. 185, New Eyes to See Inside the Sun and Stars, ed. F.-L. Deubner, J. Christensen-Dalsgaard, & D. Kurtz, 167
- Appourchaux et al. (1998c) Appourchaux, T., Rabello-Soares, M.-C., & Gizon, L. 1998c, A&AS, 132, 121
- Appourchaux & Toutain (1998) Appourchaux, T. & Toutain, T. 1998, in Sounding Solar and Stellar Interiors, IAU 181, Poster volume, ed. J. Provost & F.-X. Schmider (Kluwer Academic Publishers, Dordrecht), 5–8
- Appourchaux et al. (1995) Appourchaux, T., Toutain, T., Telljohann, U., et al. 1995, A&A, 294, L13
- Baglin (2006) Baglin, A. 2006, The CoRoT mission, pre-launch status, stellar seismology and planet finding (M.Fridlund, A.Baglin, J.Lochard and L.Conroy eds, ESA SP-1306, ESA Publication Division, Noordwijk, The Netherlands)
- Ballot et al. (2011) Ballot, J., Barban, C., & van’t Veer-Menneret, C. 2011, A&A, 531, A124
- Ballot et al. (2021) Ballot, J., Roudier, T., Malherbe, J. M., & Frank, Z. 2021, A&A, 652, A103
- Benomar et al. (2009) Benomar, O., Appourchaux, T., & Baudin, F. 2009, A&A, 506, 15
- Benomar et al. (2023) Benomar, O., Takata, M., Bazot, M., et al. 2023, A&A, 680, A27
- Broomhall et al. (2007) Broomhall, A. M., Chaplin, W. J., Elsworth, Y., & Appourchaux, T. 2007, MNRAS, 379, 2
- Broomhall et al. (2015) Broomhall, A. M., Pugh, C. E., & Nakariakov, V. M. 2015, Advances in Space Research, 56, 2706
- Carpenter et al. (2010) Carpenter, K. G., Schrijver, C. J., Karovska, M., et al. 2010, ArXiv e-prints, 1011.5214 [arXiv:1011.5214]
- Chaplin et al. (1997) Chaplin, W. J., Elsworth, Y., Isaak, G. R., et al. 1997, MNRAS, 288, 623
- Chaplin et al. (2000) Chaplin, W. J., Elsworth, Y., Isaak, G. R., Miller, B. A., & New, R. 2000, MNRAS, 313, 32
- Chaplin et al. (2004) Chaplin, W. J., Elsworth, Y., Isaak, G. R., Miller, B. A., & New, R. 2004, MNRAS, 352, 1102
- Chaplin et al. (2008) Chaplin, W. J., Houdek, G., Appourchaux, T., et al. 2008, ArXiv e-prints, 804 [0804.4371]
- Claverie et al. (1979) Claverie, A., Isaak, G., McLeod, C., van der Raay, H., & Roca Cortés, T. 1979, Nature, 282, 591
- Corbard et al. (2013) Corbard, T., Salabert, D., Boumier, P., et al. 2013, in Journal of Physics Conference Series, Vol. 440, Journal of Physics Conference Series, 012025
- Corsaro & De Ridder (2014) Corsaro, E. & De Ridder, J. 2014, A&A, 571, A71
- Davies et al. (2014) Davies, G. R., Broomhall, A. M., Chaplin, W. J., Elsworth, Y., & Hale, S. J. 2014, MNRAS, 439, 2025
- Duvall et al. (1984) Duvall, T. L., J., Dziembowski, W. A., Goode, P. R., et al. 1984, Nature, 310, 22
- Duvall & Harvey (1984) Duvall, T. L., J. & Harvey, J. W. 1984, Nature, 310, 19
- Echeverria et al. (2016) Echeverria, A., Silva, J. F., Mendez, R. A., & Orchard, M. 2016, A&A, 594, A111
- Egamberdiev & Fossat (1991) Egamberdiev, S. & Fossat, E. 1991, Sol. Phys., 133
- Elsworth et al. (1991) Elsworth, Y., Howe, R., Isaak, G. R., McLeod, C. P., & New, R. 1991, MNRAS, 251, 7P
- Fournier et al. (2025) Fournier, D., Kostogryz, N. M., Gizon, L., et al. 2025, arXiv e-prints, arXiv:2507.13756
- Fröhlich et al. (1988) Fröhlich, C., Bonnet, R. M., Bruns, A. V., et al. 1988, in Seismology of the Sun and Sun-Like Stars, ed. V.Domingo & E.Rolfe (ESA SP-286, ESA Publications Division, Noordwijk, The Netherlands), 359
- Fröhlich et al. (1995) Fröhlich, C., Romero, J., Roth, H., et al. 1995, Sol. Phys., 162, 101
- Fröhlich & Van der Raay (1984) Fröhlich, C. & Van der Raay, H. B. 1984, in The Hydromagnetics of the Sun, ed. T.D.Guyenne & J.J.Hunt (ESA SP-220, ESA Publications Division, Noordwijk , The Netherlands), 17
- Gabriel et al. (1995) Gabriel, A. H., Grec, G., Charra, J., et al. 1995, Sol. Phys., 162, 61
- Gabriel (1994) Gabriel, M. 1994, A&A, 287, 685
- García et al. (2011) García, R. A., Salabert, D., Ballot, J., et al. 2011, in Journal of Physics Conference Series, Vol. 271, GONG-SoHO 24: A New Era of Seismology of the Sun and Solar-Like Stars, 012049
- Gelly et al. (2002) Gelly, B., Lazrek, M., Grec, G., et al. 2002, A&A, 394, 285
- Gilliland et al. (2010) Gilliland, R. L., Brown, T. M., Christensen-Dalsgaard, J., et al. 2010, PASP, 122, 131
- Gizon et al. (1998) Gizon, L., Appourchaux, T., & Gough, D. O. 1998, in New eyes to see inside the sun and the stars, IAU 185, ed. F.-L. Deubner, J.Christensen-Dalsgaard, & D.Kurtz, Vol. 185 (Kluwer Academic Publishers, Dordrecht, The Netherlands), 37–40
- Gough (1984) Gough, D. O. 1984, Royal Society of London Philosophical Transactions Series, 313, 27
- Gould (1855) Gould, B. A. 1855, AJ, 4, 81
- Grec (1981) Grec, G. 1981, PhD thesis, Université de Nice
- Grec et al. (1980) Grec, G., Fossat, E., & Pomerantz, M. 1980, Nature, 288, 541
- Handberg & Campante (2011) Handberg, R. & Campante, T. L. 2011, A&A, 527, A56
- Howe et al. (2023) Howe, R., Chaplin, W. J., Elsworth, Y. P., Hale, S. J., & Nielsen, M. B. 2023, MNRAS, 526, 1447
- Jiménez-Reyes et al. (2003) Jiménez-Reyes, S. J., García, R. A., Jiménez, A., & Chaplin, W. J. 2003, ApJ, 595, 446
- Keith-Hardy et al. (2019) Keith-Hardy, J. Z., Tripathy, S. C., & Jain, K. 2019, ApJ, 877, 148
- Kiefer et al. (2018) Kiefer, R., Komm, R., Hill, F., Broomhall, A.-M., & Roth, M. 2018, Sol. Phys., 293, 151
- Komm et al. (2000a) Komm, R. W., Howe, R., & Hill, F. 2000a, ApJ, 531, 1094
- Komm et al. (2000b) Komm, R. W., Howe, R., & Hill, F. 2000b, ApJ, 543, 472
- Korzennik (2023) Korzennik, S. G. 2023, Frontiers in Astronomy and Space Sciences, 9, 1031313
- Kosovichev et al. (1992) Kosovichev, A. G., Christensen-Dalsgaard, J., Daeppen, W., et al. 1992, MNRAS, 259, 536
- Kostogryz et al. (2021) Kostogryz, N. M., Fournier, D., & Gizon, L. 2021, A&A, 654, A1
- Libbrecht & Woodard (1990) Libbrecht, K. G. & Woodard, M. F. 1990, Nature, 345, 779
- Lund et al. (2017) Lund, M. N., Silva Aguirre, V., Davies, G. R., et al. 2017, ApJ, 835, 172
- Monnier & al. (2019) Monnier, J. D. & al., . 2019, arXiv e-prints, arXiv:1907.09583
- Neckel & Labs (1994) Neckel, H. & Labs, D. 1994, Sol. Phys., 153, 91
- Neiner & Appourchaux (2004) Neiner, C. & Appourchaux, T. 2004, in ESA Special Publication, Vol. 538, Stellar Structure and Habitable Planet Finding, ed. F. Favata, S. Aigrain, & A. Wilson, 373–375
- Peirce (1852) Peirce, B. 1852, AJ, 2, 161
- Prokhorov et al. (2014) Prokhorov, A., Zhugzhda, Y. D., & Berdyugina, S. 2014, Astronomische Nachrichten, 335, 150
- Rauer et al. (2025) Rauer, H., Aerts, C., Cabrera, J., et al. 2025, Experimental Astronomy, 59, 26
- Ricker et al. (2014) Ricker, G. R., Winn, J. N., Vanderspek, R., et al. 2014, Journal of Astronomical Telescopes, Instruments, and Systems, 1, 014003
- Ritzwoller & Lavely (1991) Ritzwoller, M. H. & Lavely, E. M. 1991, ApJ, 369, 557
- Salabert et al. (2009) Salabert, D., Leibacher, J., Appourchaux, T., & Hill, F. 2009, ApJ, 696, 653
- Scherrer et al. (1995) Scherrer, P. H., Bogart, R. S., Bush, R. I., et al. 1995, Sol. Phys., 162, 129
- Scherrer et al. (2012) Scherrer, P. H., Schou, J., Bush, R. I., et al. 2012, Sol. Phys., 275, 207
- Schou (1992) Schou, J. 1992, PhD thesis, Århus Universitet, Denmark
- Schou et al. (2002) Schou, J., Howe, R., Basu, S., et al. 2002, ApJ, 567, 1234
- Siddiqi (2016) Siddiqi, A. 2016, Beyond Earth - A Chronicle of deep space exploration, 19582016, NASA SP-2018-4041, Tech. rep., National Aeronautics and Space Administration
- Stahn & Gizon (2008) Stahn, T. & Gizon, L. 2008, Sol. Phys., 251, 31
- Toner et al. (1999) Toner, C. G., Jefferies, S. M., & Toutain, T. 1999, ApJ, 518, L127
- Toutain et al. (1997) Toutain, T., Appourchaux, T., Baudin, F., et al. 1997, Sol. Phys., 175, 311
- Toutain et al. (1998) Toutain, T., Appourchaux, T., Fröhlich, C., et al. 1998, ApJ, 506, L147
- Toutain et al. (1999) Toutain, T., Berthomieu, G., & Provost, J. 1999, A&A, 344, 188
- Toutain & Fröhlich (1992) Toutain, T. & Fröhlich, C. 1992, A&A, 257, 287
- Toutain & Gouttebroze (1993) Toutain, T. & Gouttebroze, P. 1993, A&A, 268, 309
- Toutain & Kosovichev (2005) Toutain, T. & Kosovichev, A. G. 2005, ApJ, 622, 1314
- Van Trees (1968) Van Trees, H. L. 1968, Inc., New York
- Vasilyev & Gizon (2024) Vasilyev, V. & Gizon, L. 2024, A&A, 682, A142
- Willson (1979) Willson, R. C. 1979, Appl. Opt., 18, 179
- Woodard & Hudson (1983) Woodard, M. & Hudson, H. 1983, Sol. Phys., 82, 67
Appendix A Clebsh-Gordan expansion of the mode splitting
In the early days of helioseismology (before the 1990’s), mode splittings were usually expressed in term of Legendre polynomials. Unfortunately, these polynomials are orthogonal only on a continuous space (between [-1,1]) not on a discrete set such as (-1, 0,+1) for =1, for example. Therefore, other expansion are required that can either be computed by hand or derived from quantum mechanics. Ritzwoller & Lavely (1991) derived these polynomials from quantum mechanics and used a normalization such that even for high degree modes their are asymptotically equal to the usual Legendre polynomials. It is now common practice to have a normalization such that this is the case. The splittings are expressed as follows:
where is the degree, the order, and selected are given as follows:
with . Please note that for all we have . The polynomials are derived from Eqs (39) to (44) of Ritzwoller & Lavely (1991).
Appendix B LOI visibilities
The LOI visibilities are written as:
where is the number of pixels, and are given by
where is the domain representing pixel , is the spherical harmonic and are the spherical angular coordinates of a point on the surface of the Sun and is the limb darkening.
Please note that with the definition of Eq. (11), , and therefore the visibility of the LOI data for is simply .
| 0 | 0 | 1.000 |
| 1 | 0 | 0.749 |
| 1 | 1 | 1.069 |
| 2 | 0 | 0.744 |
| 2 | 1 | 0.877 |
| 2 | 2 | 1.045 |
| 3 | 0 | 0.695 |
| 3 | 1 | 0.783 |
| 3 | 2 | 0.857 |
| 3 | 3 | 0.983 |
Appendix C HV guiding signals and guiding pixel visibilities
The HV guiding signals are proportional to the difference in the guiding pixel fluxes (EW or NS). Therefore, for computing the ratio of visibilities for the HV guiding signals, one has to compute the intensity perturbation obtained from the guiding pixels.
The reference axis for spherical harmonics is linked to the solar rotation axis, being in the direction of the observer, being the solar rotation axis and making the axis system right-handed. The corresponding spherical coordinates are
Since we are trying to integrate the flux over the guiding detectors having a circular symmetry, we use the system with pointing toward the observer. A rotation by 90 degrees transforms the system ) to . The associated spherical coordinates are then . In that framework, the East and West guiding pixels are defined by [-135,-45] and [+45,+135], respectively; while the North and South guiding pixels are defined by by [-45, +45] and [+135,+225].
C.1 Intensity perturbation
Assuming that the intensity perturbation is proportional to the spherical harmonics, the perturbation integrated over the guiding pixels for a given , can be written as:
where is the projected surface element, are the spherical angular coordinates, and is the limb darkening. The spherical harmonics expressed in the system can be decomposed using Edmonds rotation matrices as a function of spherical harmonics expressed in the system as:
Since we also have the following expression :
We can rewrite Eq. (C.2) as:
Using Eq. (C.3) we can then rewrite Eq. (C.1) as:
The surface element can be written as , with . Equation (C.5) can then be separated in and as:
We can rewrite Eq. (C.6) as:
with
and:
Since we have , we have also
Here I note that the perturbation used in Eq. (C.1) is supposed to be a spherical harmonic. Toutain et al. (1999) demonstrated that any perturbation can be decomposed using Edmonds rotations matrices. As a result, Eq. (C.7) does not only apply to spherical harmonics but to any perturbation using spherical harmonics as a base, involving its derivatives, or any other combination.
C.2 for the E and W guiding pixels
The guiding pixel integrals can be calculated for the E and W pixels (E: [-135,-45] and W: [+45,+135]):
and
C.3 for the N and S guiding pixel
The guiding pixel integrals can be calculated for the N and S pixels (N: [-45, +45] and S: [+135,+225]):
and
C.4 Rotation matrices
C.5 for the HV voltage
The HV signals are proportional to the difference in the intensity between the E (N) and W (S) pixels. Then we can write:
and
Using the rotation matrices given above, Eq. (C.10) to (C.13) and Eq. (C.7), Tables (C.1) to (C.3) provides the results for Eq. (C.17) and (C.18). Table (C.4) gives the resulting visibilities in power after summing over all components in a multiplet.
| Order | EW | NS |
| 0 | ||
| 0 |
| Order | EW | NS |
| 0 | ||
| 0 | ||
| 0 | 0 |
| Order | EW | NS |
| 0 | ||
| 0 | ||
| 0 | ||
| 0 |
| Degree | EW | NS |
| 1 | 16 | 16 |
| 2 | 16 | 16 |
| 3 | 16 + 16 /9 | 16 + 16 /9 |
C.6 for the EW and NS guiding pixels
The EW guiding pixel signals are proportional to the mean intensity of the E (N) and W (S) pixels. Then we can write:
and
| Order | EW | NS |
| 0 | 0 |
| Order | EW | NS |
| 0 | 0 | |
| Degree | EW | NS |
| 1 | ||
| 2 | 5 | 5 |
C.7 The multiplet
C.7.1 =1 modes
In this case, the responses for the EW / NS pixels and the HV EW signals are different from zero, while nothing is transmitted for . For the NS HV signal, only is transmitted. The visibilities111sum over the transmitted modes in the power spectrum for the EW signals and for the NS signals are identical.
C.7.2 =2 modes
In this case, the responses for the EW guiding signals are different from zero, while nothing is transmitted for other . For the NS guiding signals, only are transmitted. For the EW and NS pixels, the responses are different from 0 for and . The visibilities in the power spectrum for the EW signals and for the NS signals are identical.
C.7.3 =3 modes
In this case, the responses for the EW guiding signals are different from zero, while nothing is transmitted for other . For the NS guiding signals, only are transmitted. The visibilities in the power spectrum for EW and NS guiding signals are identical.
Appendix D Observed visibility ratios
| Ratio | theor. | obs. | diff. |
| +0.488 | +0.498 0.030 | +0.010 0.020 | |
| -0.446 | -0.444 0.009 | +0.002 0.009 | |
| -3.912 | -3.276 0.028 | +0.636 0.020 |
| Degree | Ratio | theor. | obs. | diff. |
| 1 | -0.701 | -0.684 0.078 | +0.017 0.078 | |
| 2 | -0.674 | -0.464 0.046 | +0.210 0.046 | |
| 3 | -0.348 | -0.286 0.038 | +0.062 0.038 | |
| 3 | -0.698 | -0.677 0.066 | +0.021 0.066 | |
| 3 | -0.448 | -0.489 0.052 | –0.041 0.052 | |
| 3 | -0.274 | -0.341 0.034 | –0.067 0.034 | |
| 1 | / FD | -0.068 | -0.266 0.020 | –0.198 0.020 |
| 2 | / FD | +0.821 | 0.653 0.019 | –0.180 0.019 |
| Ratio | theor. | obs. | diff. |
| (EW) | +0.762 | +0.717 0.030 | -0.045 0.020 |
| (EW) | +0.216 | +0.340 0.008 | +0.124 0.008 |
| (NS) | +0.730 | +0.716 0.020 | -0.014 0.020 |
| (NS) | +0.318 | +0.234 0.008 | -0.084 0.007 |
| EW / FD | -0.652 | -0.685 0.061 | -0.023 0.060 |
| EW / FD | +1.059 | +0.982 0.017 | -0.077 0.060 |
| NS / FD | -0.770 | -0.707 0.063 | +0.063 0.060 |
| NS / FD | +0.909 | +0.934 0.022 | +0.025 0.060 |
| Ratio | theor. | obs. | diff. |
| (EWHV) | -1.570 | 0.848 0.042 | +2.418 0.042 |
| (EWHV) | -0.414 | 0.455 0.028 | +0.869 0.028 |
| (NSHV) | -1.570 | 0.762 0.047 | +2.332 0.047 |
| (NSHV) | -0.414 | 0.539 0.036 | +0.953 0.036 |
| (EW / NS)HV | 0. | 0.349 0.067 | +0.349 0.067 |
| (EW / NS)HV | 0. | 0.411 0.039 | +0.411 0.039 |
| Ratio | theor. | obs. | diff. |
| EW HV/gui. | +2.91 | 0.22 | -2.70 |
| NS HV/gui. | +2.91 | 1.61 | -1.30 |
| EW HV/gui. | -1.56 | 0.07 | +1.63 |
| NS HV/gui. | -1.56 | 0.91 | +2.57 |
Appendix E Frequencies
| 11 | 1686.599 0.011 | - | - | - |
| 12 | 1822.300 0.044 | 1885.134 0.023 | 1945.860 0.057 | |
| 13 | 1957.538 0.033 | 2020.900 0.030 | 2082.156 0.036 | |
| 14 | 2093.543 0.019 | 2156.918 0.025 | 2217.730 0.048 | |
| 15 | 2228.846 0.021 | 2292.082 0.015 | 2352.384 0.042 | |
| 16 | 2362.889 0.020 | 2425.705 0.013 | 2486.010 0.030 | |
| 17 | 2496.286 0.017 | 2559.291 0.015 | 2619.766 0.020 | |
| 18 | 2629.754 0.010 | 2693.494 0.009 | 2754.594 0.015 | 2811.334 0.079 |
| 19 | 2764.271 0.011 | 2828.281 0.012 | 2889.676 0.015 | 2946.893 0.048 |
| 20 | 2899.098 0.012 | 2963.441 0.012 | 3024.819 0.013 | 3082.315 0.032 |
| 21 | 3033.864 0.012 | 3098.310 0.011 | 3159.989 0.016 | 3217.731 0.038 |
| 22 | 3168.738 0.013 | 3233.317 0.013 | 3295.155 0.020 | 3353.318 0.067 |
| 23 | 3303.663 0.016 | 3368.741 0.016 | 3430.861 0.028 | 3489.452 0.104 |
| 24 | 3439.152 0.023 | 3504.460 0.023 | 3566.921 0.043 | 3625.931 0.179 |
| 25 | 3575.007 0.028 | 3640.773 0.032 | 3703.428 0.067 | 3762.616 0.360 |
| 26 | 3711.196 0.051 | 3777.291 0.044 | 3840.297 0.096 | 3900.347 0.595 |
| 27 | 3847.985 0.057 | 3914.350 0.064 | 3977.802 0.189 | 4043.565 0.016 |
| 28 | 3985.001 0.122 | 4051.804 0.084 | 4115.178 0.199 | 4197.790 0.893 |
| 29 | 4122.931 0.147 | 4189.064 0.107 | 4262.748 0.087 | 4339.006 0.900 |
| 30 | 4255.467 0.075 | 4327.551 0.118 | 4391.455 0.155 | 4473.272 1.319 |
| 31 | 4399.915 0.192 | 4466.338 0.175 | 4530.130 0.178 | |
| 32 | 4538.289 0.140 | 4605.468 0.237 | 4683.535 0.417 | |
| 33 | 4672.324 0.168 | 4745.060 0.261 | - |
| 9 | - | - | - | 1591.559 0.012 |
| 10 | - | - | 1674.546 0.015 | 1729.144 0.021 |
| 11 | - | 1749.334 0.028 | 1810.301 0.016 | - |
| 12 | - | 1885.145 0.020 | 1945.886 0.016 | - |
Appendix F Linewidths and mode heights
| frequency (Hz) | log(linewidth) | log(height) | |
| 11 | 1674.55 | -1.688 0.333 | - |
| 12 | 1810.30 | -1.181 0.191 | - |
| 13 | 1957.54 | -1.666 0.084 | -0.785 0.097 |
| 14 | 2093.54 | -1.017 0.046 | -0.820 0.048 |
| 15 | 2228.85 | -0.995 0.032 | -0.340 0.032 |
| 16 | 2362.89 | -0.641 0.025 | -0.149 0.023 |
| 17 | 2496.29 | -0.228 0.020 | -0.022 0.017 |
| 18 | 2629.75 | +0.309 0.026 | -0.085 0.019 |
| 19 | 2764.27 | +0.711 0.015 | -0.052 0.010 |
| 20 | 2899.10 | +1.079 0.012 | -0.096 0.008 |
| 21 | 3033.86 | +1.230 0.013 | -0.015 0.008 |
| 22 | 3168.74 | +0.993 0.010 | +0.227 0.007 |
| 23 | 3303.66 | +0.512 0.011 | +0.508 0.007 |
| 24 | 3439.15 | -0.281 0.011 | +0.936 0.009 |
| 25 | 3575.01 | -1.055 0.016 | +1.306 0.012 |
| 26 | 3711.20 | -1.782 0.009 | +1.661 0.009 |
| 27 | 3847.98 | -2.493 0.015 | +1.974 0.012 |
| 28 | 3985.00 | -3.087 0.013 | +2.241 0.013 |
| 29 | 4122.93 | -3.625 0.016 | +2.408 0.014 |
| 30 | 4255.47 | -4.121 0.017 | +2.613 0.014 |
| 31 | 4399.91 | -4.509 0.026 | +2.639 0.028 |
| 32 | 4538.29 | -5.063 0.025 | +2.668 0.022 |
| 33 | 4672.32 | -5.260 0.024 | +2.836 0.023 |
| frequency (Hz) | log(linewidth) | log(height) | |
| 10 | 1591.56 | -2.492 0.566 | - |
| 11 | 1729.14 | -1.565 0.246 | - |
| 12 | 1885.13 | -1.328 0.132 | -1.041 0.148 |
| 13 | 2020.90 | -0.811 0.070 | -0.477 0.074 |
| 14 | 2156.92 | -0.572 0.040 | -0.494 0.041 |
| 15 | 2292.08 | -0.149 0.044 | -0.413 0.038 |
| 16 | 2425.70 | +0.157 0.041 | -0.137 0.033 |
| 17 | 2559.29 | +0.521 0.018 | -0.056 0.013 |
| 18 | 2693.49 | +0.935 0.020 | -0.056 0.013 |
| 19 | 2828.28 | +1.319 0.014 | -0.065 0.009 |
| 20 | 2963.44 | +1.573 0.014 | -0.020 0.009 |
| 21 | 3098.31 | +1.579 0.013 | +0.105 0.008 |
| 22 | 3233.32 | +1.272 0.015 | +0.331 0.009 |
| 23 | 3368.74 | +0.674 0.012 | +0.654 0.008 |
| 24 | 3504.46 | -0.086 0.010 | +1.038 0.007 |
| 25 | 3640.77 | -0.882 0.010 | +1.418 0.007 |
| 26 | 3777.29 | -1.597 0.010 | +1.732 0.008 |
| 27 | 3914.35 | -2.285 0.011 | +2.054 0.010 |
| 28 | 4051.80 | -2.854 0.018 | +2.286 0.016 |
| 29 | 4189.06 | -3.439 0.017 | +2.506 0.014 |
| 30 | 4327.55 | -3.895 0.020 | +2.623 0.019 |
| 31 | 4466.34 | -4.391 0.028 | +2.812 0.030 |
| 32 | 4605.47 | -4.764 0.029 | +2.910 0.034 |
| 33 | 4745.06 | -5.169 0.035 | +2.876 0.052 |

### v2

@@
 (1979), confirmed by Grec et al.
-(1980), using radial velocity measurements; and by Woodard & Hudson (1983) using irradiance measurements with ACRIM111Active Cavity Radiometer Irradiance Monitor (Willson, 1979) onboard SMM222Solar Maximum Mission.
+(1980), using radial velocity measurements; and by Woodard & Hudson (1983) using irradiance measurements with ACRIM Monitor (Active Cavity Radiometer Irradiance, Willson, 1979) onboard SMM (Solar Maximum Mission).
 These breakthroughs were followed by the first inversion of the solar structure performed by Gough (1984) and Duvall et al.
@@
 Helioseismology really became a useful tool when it demonstrated that the equation of state could be in error, and was indeed in error (Kosovichev et al., 1992).
-The diagnostic potential of helioseismology for our physic had already been perceived in the mid 80’s when the space mission SoHO333Solar and Heliospheric Observatory, an ESA/NASA mission (Special volumes of Solar Physics, 1995, 162;1997, 170; et 1997,175) was approved.
-The need for uninterrupted data can be fulfilled by a space mission, but also by ground-based networks of instrument, networks on which the Sun never sets: GONG444Global Oscillation Network Group (dedicated volume of Science, 1996, 272), IRIS555Investigation on the Rotation and Interior of the Sun (Egamberdiev & Fossat, 1991) and BiSON666Birmingham Solar-Oscillations Network (Elsworth et al., 1991).
+The diagnostic potential of helioseismology for our physic had already been perceived in the mid 80’s when the space mission SoHO (Solar and Heliospheric Observatory, an ESA/NASA mission, Special volumes of Solar Physics, 1995, vol.
+162;1997, vol.
+170; and 1997,vol.
+175)) was approved.
+The need for uninterrupted data can be fulfilled by a space mission, but also by ground-based networks of instrument, networks on which the Sun never sets: GONG (Global Oscillation Network Group, dedicated volume of Science, 1996, vol.
+272), IRIS (Investigation on the Rotation and Interior of the Sun, Egamberdiev & Fossat, 1991) and BiSON (Birmingham Solar-Oscillations Network, Elsworth et al., 1991).
 The detection of solar oscillations using solar radial velocities is feasible from the ground.
@@
 Following the detection of p modes with the ACRIM instrument, Fröhlich & Van der Raay (1984) used photometers aboard balloons to reduce the impact of the Earth atmosphere on the measurement.
-A similar photometer (IPHIR777InterPlanetary Helioseismology by IRradiance measurements, (Fröhlich et al., 1988)) was flown aboard the PHOBOS II888Soviet Mars exploration mission, (See Siddiqi, 2016) mission.
+A similar photometer (IPHIR: InterPlanetary Helioseismology by IRradiance measurements, Fröhlich et al., 1988) was flown aboard PHOBOS II (Soviet Mars exploration mission, see Siddiqi, 2016).
 This instrument returned excellent helioseismic measurements as shown by Toutain & Fröhlich (1992).
@@
 Following the at-the-times recent developments in helioseismology, three helioseismic instruments were included for obtaining solar radial-velocity and irradiance measurements using either imagers or looking at the Sun as a star.
-GOLF999Global Oscillations at Low Frequency (Gabriel et al., 1995) uses a Sodium cell to measure solar radial velocities look at the Sun as a star, while MDI 101010Michelson Doppler Imager (Scherrer et al., 1995) uses a one-million pixel detector for obtaining spatially resolved data.
-The VIRGO111111Variability of IRradiance and Gravity Oscillations (Fröhlich et al., 1995) instrument measures the irradiance of the Sun as a star, the spectral radiance with photometers, and the spectral radiance spatially resolved over the solar disk.
+GOLF (Global Oscillations at Low Frequency, Gabriel et al., 1995) uses a Sodium cell to measure solar radial velocities look at the Sun as a star, while MDI (Michelson Doppler Imager, Scherrer et al., 1995) uses a one-million pixel detector for obtaining spatially resolved data.
+The VIRGO (Variability of IRradiance and Gravity Oscillations, Fröhlich et al., 1995) instrument measures the irradiance of the Sun as a star, the spectral radiance with photometers, and the spectral radiance spatially resolved over the solar disk.
 The latter is performed by the Luminosity Oscillations Imager (LOI) whose concept was described by Andersen et al.
@@
 Its design was specifically aiming at detecting modes with a different sensitivity for degree greater than 1; thereby optimizing the chance of detecting the solar g modes using different signals.
-Instruments that detects p modes in velocity, such as GOLF and BiSON (Broomhall et al., 2007) or HMI121212Helioseismic and Magnetic Imager (Scherrer et al., 2012) and GONG (Appourchaux, 2024), benefit from a better signal-to-noise ratio allowing them to detect modes below 1 Hz.
+Instruments that detects p modes in velocity, such as GOLF and BiSON (Broomhall et al., 2007) or HMI (Helioseismic and Magnetic Imager, Scherrer et al., 2012) and GONG (Appourchaux, 2024), benefit from a better signal-to-noise ratio allowing them to detect modes below 1 Hz.
 In addition, instruments such as GONG, MDI and HMI make images of the Sun permitting them to derive by inversion the internal structure from the surface below the convection zone.
@@
 After May 25, 2003, the HGA failed to move for sending data towards the Earth.
-It was then decided to leave the HGA in a fixed position and rotate the spacecraft by 180 degrees every 3 months for transmitting to Earth131313See https://soho.nascom.nasa.gov/soc/soho_events/SOHO-Spacecraft-Events.csv.
+It was then decided to leave the HGA in a fixed position and rotate the spacecraft by 180 degrees every 3 months for transmitting to Earth111See https://soho.nascom.nasa.gov/soc/soho_events/SOHO-Spacecraft-Events.csv.
 When the S/C is rolled by 180 degrees, the pixel numbering is redefined in the process, i.e.
@@
 Therefore, this average minimizes the effect of the angle on the data.
-After combining the pixels with the filters, the time series are Fourier transformed; the positive frequencies provide the signal for , and the negative frequencies provide the conjugate of the signal for Appourchaux et al.
-(1998a).
+After combining the pixels with the filters, the time series are Fourier transformed; the positive frequencies provide the signal for , and the negative frequencies provide the conjugate of the signal for (Appourchaux et al., 1998a).
 For each harmonic, I extracted two data sets: one time series of 29 years duration and 29 time series of one-year duration.
@@
 The code was then made fully functional and was tested for the PLATO mission on hundreds of solar-like stars.
-The code can be found on Gitlab141414git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
+The code can be found on Gitlab222git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
 /msap3_04/progfit.
@@
 Since these signals are by construction not sensitive to =0 modes, the new code (guiFIT) took into account this lack of visibility, but also the different visibilities in a multiplet.
-The code can also be found on Gitlab151515git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
+The code can also be found on Gitlab333git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
 /msap3_04/guifit.
@@
 The 10.7-cm flux could be downloaded at the SIDC (Solar Influence Data Analysis Center) in 2023 with a time series running from April 1st 1996 to March 31st, 2021; the data are at the time of writing no longer available.
-The flux can now be found daily thanks to Space Weather Canada161616spaceweather.gc.ca/forecast-prevision/solar-solaire/solarflux/sx-5-en.php, and the two time series were merged into a single time series from April 1st 1996 to March 31st, 2025, for producing a yearly average.
+The flux can now be found daily thanks to Space Weather Canada444spaceweather.gc.ca/forecast-prevision/solar-solaire/solarflux/sx-5-en.php, and the two time series were merged into a single time series from April 1st 1996 to March 31st, 2025, for producing a yearly average.
 Since I used the 29 yearly time series of the different signals for that study, I also averaged the flux over the same period of one year.
@@
 9.1 Limb darkening and solar radius calibration
-The computation of the mode visibilities requires calibrating the limb darkening and the effective size of the solar radius on the LOI detector (See Appendix B).
+The computation of the mode visibilities requires calibrating the limb darkening and the effective size of the solar radius on the LOI detector (See Appendix refAppendixB).
 The calibration was done by computing the ratio of the mean of the NS pixels (1, 2.
 3, 4) to the Central pixels (9, 10, 11, 12), the ratio of the mean of the EW pixels (5, 6, 7, 8) to the Central pixels, and the ratio of the mean of the guiding pixels (12, 13, 14, 15) as a function of the SoHO-Sun distance in AU.
-The computation of the relative fluxes was performed using the shape of the pixels and the presence of the tracks as provided by the manufacturer171717Advanced Micro Electronics, Horten, Norway.
+The computation of the relative fluxes was performed using the shape of the pixels and the presence of the tracks as provided by the manufacturer555Advanced Micro Electronics, Horten, Norway.
 The tracks are non sensitive part of the detector required for getting the current resulting from the photoelectric effect.
@@
 (2013) with MDI, HMI and the PICARD mission.
-On the other hand, the visibility ratios for the EW and NS signals (guiding pixels and HV) reported in this article are clearly not in agreement with the expectation of the simpleI theory provided in Appendix C.
-There are several discrepancies with theory to be noted (See Table D.4 and D.5): 1) the mode visibilities of the EW signals are systematically higher than those of the NS signals; 2) the mode visibilities of the modes are significantly higher than those of the modes for the guiding pixels and the HV signals.
-If one compares the results of the HV signals with the EW and NS difference (See Table D;3), it is clear that for these latter signals 1) the mode visibilities in EW and NS are identical for and ; and 2) the ratio of mode visibilities of the to the modes are in agreement with theory.
+On the other hand, the visibility ratios for the EW and NS signals (guiding pixels and HV) reported in this article are clearly not in agreement with the expectation of the simple theory provided in Appendix C.
+There are several discrepancies with theory to be noted (See Table 14 and 15): 1) the mode visibilities of the EW signals are systematically higher than those of the NS signals; 2) the mode visibilities of the modes are significantly higher than those of the modes for the guiding pixels and the HV signals.
+If one compares the results of the HV signals with the EW and NS difference (See Table 13), it is clear that for these latter signals 1) the mode visibilities in EW and NS are identical for and ; and 2) the ratio of mode visibilities of the to the modes are in agreement with theory.
 It means that the intensity perturbation at the center of the disk () down to follows a product of the limb darkening with a spherical harmonic (Eq.
@@
 I show in Appendix C that the ratio of these visibilities for either the EW or NS signals is .
-For spherical harmonics, this ratio is about 0.2 (See Table D.4).
+For spherical harmonics perturbations, this ratio is about 0.2 (See Table 14).
 Using the relative intensity perturbation provided by Toutain et al.
@@
 The intensity perturbations provided by Toutain et al.
-(1999) do not contradict the simple theory provided by Table D.4.
+(1999) do not contradict the simple theory provided by Table 14.
 The theory of Toutain et al.
@@
 - Elsworth et al. (1991) Elsworth, Y., Howe, R., Isaak, G. R., McLeod, C. P., & New, R. 1991, MNRAS, 251, 7P
-- Fournier et al. (2025) Fournier, D., Kostogryz, N. M., Gizon, L., et al. 2025, arXiv e-prints, arXiv:2507.13756
+- Fournier et al. (2025) Fournier, D., Kostogryz, N., Gizon, L., et al. 2025, A&A, 702, A253
 - Fröhlich et al.
@@
 For the NS HV signal, only is transmitted.
-The visibilities111sum over the transmitted modes in the power spectrum for the EW signals and for the NS signals are identical.
+The visibilities666sum over the transmitted modes in the power spectrum for the EW signals and for the NS signals are identical.
 C.7.2 =2 modes
@@
 Appendix E Frequencies
-| 11 | 1686.599 0.011 | - | - | - |
-| 12 | 1822.300 0.044 | 1885.134 0.023 | 1945.860 0.057 | |
+| 11 | 1686.599 0.011† | - | - | - |
+| 12 | 1822.300 0.044† | 1885.134 0.023 | 1945.860 0.057 | |
 | 13 | 1957.538 0.033 | 2020.900 0.030 | 2082.156 0.036 | |
@@
 | 11 | - | 1749.334 0.028 | 1810.301 0.016 | - |
-| 12 | - | 1885.145 0.020 | 1945.886 0.016 | - |
+| 12 | - | 1885.145 0.020† | 1945.886 0.016† | - |
 Appendix F Linewidths and mode heights
 | frequency (Hz) | log(linewidth) | log(height) | |
-| 11 | 1674.55 | -1.688 0.333 | - |
-| 12 | 1810.30 | -1.181 0.191 | - |
+| 11† | 1674.55 | -1.688 0.333 | - |
+| 12† | 1810.30 | -1.181 0.191 | - |
 | 13 | 1957.54 | -1.666 0.084 | -0.785 0.097 |
@@
 | frequency (Hz) | log(linewidth) | log(height) | |
-| 10 | 1591.56 | -2.492 0.566 | - |
-| 11 | 1729.14 | -1.565 0.246 | - |
+| 10† | 1591.56 | 2.492 0.566 | - |
+| 11† | 1729.14 | -1.565 0.246 | - |
 | 12 | 1885.13 | -1.328 0.132 | -1.041 0.148 |